a lens of focal length 5.0 cm is being used by a student in the laboratory as a magnifying glass. His least distance vision is 25cm. What magnification is the student getting?

Grade:10

1 Answers

Lakshit
7 years ago
we know that,
                     1/v – 1/u = 1/f
1/v = 4/25
v = 25/4
Magnification = v/u
                    = (25/4)/(-25)
                    = -1/4,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,